The average price to earnings ratio for the soft beverage industry is 25. If Coca Cola has an earnings per share of 2.10, what would you pay for this stock? DO NOT USE DOLLAR SIGNS OR COMMAS IN YOUR ANSWER. ROUND ANSWER TO THE NEAREST CENT (2 Decimals).

Your Answer:
Answer #1

Price of coca cola = P/E ratio*earning per share= 25*2.1=52.5